# LIT-LVM: Structured Regularization for Interaction Terms in Linear Predictors using Latent Variable Models

### v1

@@ Results
The regression and classification results are shown in Tables 1 and 2, respectively. Our LIT-LVM approach achieves superior accuracy over the suite of datasets for both tasks, and by a large margin on some datasets, such as nyc_taxi_green for regression and clean_2 for classification. As we observed in the simulation experiments, our proposed LIT-LVM approach achieves the highest gains compared to elastic net (with interactions) on datasets with high $p^{2}/n$ ratios, including superconduct for regression and fri_c4_1000_100 for classification. This also agrees with well-known results for other regularization methods such as elastic net, where the benefit is mainly achieved when the ratio $p/n$ approaches or exceeds 1. With the inclusion of $\binom{p}{2}$ interaction terms, it is the ratio $p^{2}/n$ that is important. We also note that, without the structured regularization of LIT-LVM, interaction terms can often hurt the accuracy of a linear predictor, as shown by classification datasets such as hill-valley, where elastic net (no interactions) far exceeds the accuracy with interactions.
@@
Figure 4: Mean AUC for different values of the LVM regularization strength $\lambda_{l}$ on real OpenML classification datasets. (4(a)) The clean1 dataset peaks in AUC at a moderate $\lambda_{l}$ between $0.01$ and $0.1$, while the more difficult fri_c4_500_100 dataset benefits from stronger LVM regularization with $\lambda_{l}$ in the range of $1$ to $100$.
The effects of $\lambda_{l}$ on mean AUC in the clean1 dataset are shown in Figure 4(a). From Table 2, we see that LIT-LVM already attains a substantially higher AUC on clean1 than the other baselines. We observe that moderate values of $\lambda_{l}$ again yield the best result, reinforcing that some, but not excessive, emphasis on latent structure can exploit the potential low-rank signal in the data. The performance on another data set, fri_c4_500_100 is shown in Figure 4(b). From Table 2, we see that the mean AUC of FM and LIT-LVM are closer on this dataset than on clean1, which indicates that LIT-LVM is indeed imposing a low-rank structure similar to FM, but less rigid, and thus, still manages to outperform it. Consequently, we observe a higher value for $\lambda_{l}$ to maximize AUC compared to clean1. Nevertheless, the overall pattern remains similar: a moderate $\lambda_{l}$ still achieves or exceeds the best possible performance from an FM, which is a purely factorized method, or a model with just an elastic net penalty and no factorization.
@@
Table 3 shows that LIT-LVM outperforms the all other variants of Cox PH models in discrimination and even matches the much more complex random-survival forest (RSF) on the C-index ($0.630$ vs. $0.629$), yet achieves the best calibration with the lowest integrated Brier score of $0.143$. The discordance for RSF, having the second-best C-index but worst IBS $(0.155)$ is not surprising, as the C-index measures ranking only, whereas the Brier score also measures calibration, and tree ensembles such as RSF have been found to suffer from poorer calibration than Cox PH models in other work Steele et al. (2018). Notably, the $\approx 0.006$ ($\approx 4\%$) reduction in IBS that LIT-LVM achieves over the elastic net Cox is an order of magnitude larger than the $\leq 0.001$ differences among the other Cox PH variants, underscoring a clear calibration advantage. The difference in IBS between LIT-LVM and the other models is further explained in Section C.2.2 by examining the Brier score over time. LIT-LVM is the only model that combines superior discrimination with superior calibration.
@@ 6 Conclusion
In this paper, we limited our investigation to second-order interaction terms. Some datasets may exhibit higher-order interactions; our proposed approach could potentially be extended to incorporate these, with the interaction coefficient matrix $\mathbf{\bm{\Theta}}$ replaced by a higher-order tensor. The approximate low-dimensional structure could then be enforced using tensor factorization approaches Kuleshov et al. (2015). Our LIT-LVM approach also excludes terms of the form $x_{j}^{2}$, which can be viewed as self interactions and are usually included in polynomial features. These self interactions could be added into LIT-LVM as terms on the diagonal of $\mathbf{\bm{\Theta}}$ and by choosing a different type of low-dimensional structure that incorporates these terms.
@@ Acknowledgments
This material is based upon work supported by the National Science Foundation grant IIS-2318751.
Research reported in this publication was supported by the National Library of Medicine of the National Institutes of Health under Award Number R01LM013311 as part of the NSF/NLM Generalizable Data Science Methods for Biomedical Research Program. The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health.
@@ A.1 Targeted Interactions
Inclusion of the interaction terms between all the features might not be well suited given the conditions and the domain pertinent to a specific application. Instead, we may be interested in incorporating the interaction between a first set of features $\mathbf{\bm{X}}^{{}^{\prime}}\subseteq\mathbf{\bm{X}}$ with dimension of $r$ and a second set of features $\mathbf{\bm{X}}^{{}^{\prime\prime}}\subseteq\mathbf{\bm{X}}$ with dimension of $q$. We refer to these as targeted interactions, which may come from domain knowledge or specific interest in a set of interactions. With targeted interactions, the interaction matrix $\mathbf{\Theta}$ changes from a square $p\times p$ to a possibly rectangular ${r\times q}$ matrix represent the coefficients for the interaction terms between the features of $\mathbf{\bm{X}}^{{}^{\prime}},\mathbf{\bm{X}}^{{}^{\prime\prime}}$. We denote the latent representation matrices of the two sets of features $\mathbf{\bm{X}}^{{}^{\prime}},\mathbf{\bm{X}}^{{}^{\prime\prime}}$ as $\mathbf{\bm{Z}}^{{}^{\prime}}_{r\times d},\mathbf{\bm{Z}}^{{}^{\prime\prime}}_{q\times d}$, respectively.
@@ B.1 Linear Regression with Low Rank Model
Figure 6: Results of simulation experiments on linear regression for different values of the true latent dimension $d_{true}$. Lower RMSE denotes higher prediction accuracy. Results for $d_{\text{true}}=2$ are shown in Figure 2(a). Error bars denote one standard error.
@@
Figure 7: Results of simulation experiments on linear regression for different values of the chosen latent dimension $d$ and true latent dimension $d_{true}$. Lower RMSE denotes higher prediction accuracy. Error bars denote one standard error.
B.2 Logistic Regression with Latent Distance Model
@@
Figure 8: Results of simulation experiments on logistic regression for different values of the true latent dimension $d_{true}$. Higher AUC denotes higher prediction accuracy. Results for $d_{\text{true}}=2$ are shown in Figure 2(b). Error bars denote one standard error.
@@
Finally, we explore what happens when we let $p$ grow to an extremely high value so that $\binom{p}{2}\gg n$. The results up to $p=500$, for which $\binom{p}{2}=124,\!750$ with $n=1,\!000$ are shown in Figure 10. Notice that LIT-LVM continues to outperform elastic net until the point where all of the methods are barely better than a random guess at $p=500$.
B.3 LIT-LVM vs. Factorization Machines (FMs)
@@
In this simulation study, we evaluate the performance of FMs compared to LIT-LVM through a series of classification experiments under varied conditions such as increased noise and sparsity injected in the interaction matrix $\mathbf{\bm{\Theta}}$. These conditions are critical for assessing the performance of the models, as they can potentially affect the predictive accuracy of models like FMs that rely on exact low-rank structures. By introducing greater noise variance and higher sparsity levels in the interaction weights during the simulations, we aim to demonstrate how each model handles deviations from the exact low-rank assumption, thereby providing insights into their practical use in scenarios where this assumption does not hold. Since we are interested in the accuracy of the FM and not the computation time, we formulate FMs as a special case of LIT-LVM with extremely large $\lambda_{l}$. (The model recovered by LIT-LVM is equivalent to that of an FM in the limit $\lambda_{l}\rightarrow\infty$ as discussed in Section 3.3.)
Figure 11 displays the mean AUC results for both models under baseline noise and sparsity conditions, while Figure 12 and 13 show the results when the models are subjected to increased noise levels and greater sparsity. Figure 12 shows the effect of increasing the noise variance to $4$ on the performance of the models, highlighting a significant drop in the accuracy of FMs due to their reliance on an exact low-rank structure of the interaction weight matrix. Similarly, Figure 13 demonstrates how increasing the sparsity level in the interaction weights impacts the performance of FMs compared to LIT-LVM. These experiments underscore the robustness of LIT-LVM under more challenging conditions where assumptions of the exact low-rank structures in FMs do not hold.
@@ C.1 OpenML Datasets
In our experiments, we utilized a diverse set of benchmark datasets to evaluate the performance of the proposed model. These datasets were chosen to ensure comprehensive coverage of different types of data distributions and characteristics. By using these datasets, we aimed to thoroughly validate the performance and generalization ability of our model across both classification and regression tasks.
@@ C.1.1 Data Description
Table 4: Regression datasets from OpenML used for real data experiments, comprising a wide range of number of samples $n$ and features $p$. Datasets with a slash in the number of features $p$ have categorical variables, so the entry denotes before/after one-hot encoding. The OpenML dataset ID column includes a link to the dataset.
Dataset Name # of Samples $n$ of Features $p$ OpenML Dataset ID
boston 506 14/22 531
Bike_Sharing_Demand 1,737 6 44142
diamonds 53,940 10 44140
fri_c4_500_10 500 10 604
medical_charges 163,065 12 44146
nyc-taxi-green-dec-2016 581,835 9 44143
pol 15,000 26 44134
rmftsa_ladata 508 10 666
socmob 1,156 6/39 541
space_ga 3,106 7 507
superconduct 21,263 79 44148
wind 6,574 15 503
@@
Dataset Name # of Samples $n$ of Features $p$ OpenML Dataset ID
Bioresponse 3,434 420 45019
clean1 476 169 40665
clean2 6598 169 40666
eye_movements 7,698 24 44157
fri_c4_500_100 500 101 742
fri_c4_1000_100 1,000 101 718
hill-valley 1,212 101 1479
jannis 57,580 55 43977
tecator 240 125 851
We employ a range of datasets as detailed in Table 4 for regression and Table 5 for classification. These datasets were selected to test the model’s ability to handle different scenarios. The diversity in the number of samples and features ensures that our model is evaluated on a wide spectrum of regression and classification challenges. Links to these datasets are also provided for ease of access.
C.1.2 Comparison of LIT-LVM with FMs
@@
C.2 Survival Prediction and Compatibility Estimation for Kidney Transplantation
C.2.1 Data Preprocessing

### v2

@@ Results
The regression and classification results are shown in Tables 1 and 2, respectively. Our LIT-LVM approach achieves superior accuracy over the suite of datasets for both tasks, and by a large margin on some datasets, such as nyc_taxi_green for regression and clean_2 for classification. As we observed in the simulation experiments, our proposed LIT-LVM approach achieves the highest gains compared to elastic net with interactions on datasets with high $p^{2}/n$ ratios (shown in Tables 4 and 5 in Section C.1.1), including superconduct for regression and fri_c4_1000_100 for classification. This also agrees with well-known results for other regularization methods such as elastic net, where the benefit is mainly achieved when the ratio $p/n$ approaches or exceeds 1. With the inclusion of $\binom{p}{2}$ interaction terms, it is the ratio $p^{2}/n$ that is important. We also note that, without the structured regularization of LIT-LVM, interaction terms can often hurt the accuracy of a linear predictor, as shown by classification datasets such as hill-valley, where elastic net without interactions far exceeds the accuracy with interactions. Such results were also observed by Bien et al. (2013) in their experiments comparing elastic net without interactions, with interactions, and hierarchical lasso. We find that the hierarchical lasso’s prediction accuracy is generally in between that of elastic net with and without interactions, as one might expect. The weak heredity assumption provides a stronger regularization compared to elastic net with interactions, which allows hierarchical lasso to perform better, but not as well as LIT-LVM, on datasets with high $p^{2}/n$ ratios. On all datasets, we find that our proposed optimization procedure for LIT-LVM converges, as shown in Section C.2.
@@
Figure 4: Mean AUC for different values of the LVM regularization strength $\lambda_{l}$ on real OpenML classification datasets. (4(a)) The clean1 dataset peaks in AUC at a moderate $\lambda_{l}$ between $0.01$ and $0.1$, while the more difficult fri_c4_500_100 dataset benefits from stronger LVM regularization with $\lambda_{l}$ in the range of 11 to $100$.
The effects of $\lambda_{l}$ on mean AUC in the clean1 dataset are shown in Figure 4(a). From Table 2, we see that LIT-LVM already attains a substantially higher AUC on clean1 than the other baselines. We observe that moderate values of $\lambda_{l}$ again yield the best result, reinforcing that some, but not excessive, emphasis on latent structure can exploit the potential low-rank signal in the data. The performance on another data set, fri_c4_500_100 is shown in Figure 4(b). From Table 2, we see that the mean AUC of FM and LIT‑LVM are closer on this dataset than on clean1, which indicates that LIT‑LVM is indeed imposing a low‑rank structure similar to FM, but less rigid, and thus, still manages to outperform it. Consequently, we observe a higher value for $\lambda_{l}$ to maximize AUC compared to clean1. Nevertheless, the overall pattern remains similar: a moderate $\lambda_{l}$ still achieves or exceeds the best possible performance from an FM, which is a purely factorized method, or a model with just an elastic net penalty and no factorization.
@@
Table 3 shows that LIT‑LVM outperforms the all other variants of Cox PH models in discrimination and even the much more complex random survival forest (RSF) on the C‑index ($0.630$ vs. $0.629$), yet also achieves the best calibration with the lowest integrated Brier score of $0.143$. The discordance for RSF, having the second‑best C‑index but worst IBS $(0.156)$ is not surprising, as the C‑index measures ranking only, whereas the Brier score also measures calibration, and tree ensembles such as RSF have been found to suffer from poorer calibration than Cox PH models in other work Steele et al. (2018). Notably, the $\approx 0.006$ ($\approx 4\%$) reduction in IBS that LIT‑LVM achieves over the elastic net Cox is an order of magnitude larger than the $\leq 0.001$ differences among the other Cox PH variants, underscoring a clear calibration advantage. The difference in IBS between LIT-LVM and the other models is further explained in Section C.4.2 by examining the Brier score over time. LIT‑LVM is the only model that combines superior discrimination with superior calibration.
@@ 6 Conclusion
In this paper, we limited our investigation to second-order interaction terms. Some datasets may exhibit higher-order interactions; our proposed approach could potentially be extended to incorporate these, with the interaction coefficient matrix $\mathbf{\bm{\Theta}}$ replaced by a higher-order tensor. The approximate low-dimensional structure could then be enforced using tensor factorization approaches Kuleshov et al. (2015). Our LIT-LVM approach also excludes terms of the form $x_{j}^{2}$, which can be viewed as self interactions and are usually included in polynomial features. These self interactions could be added into LIT-LVM as terms on the diagonal of $\mathbf{\bm{\Theta}}$ and by choosing a different type of low-dimensional structure that incorporates these terms. Our proposed approach also only scales to thousands of features, as it requires estimation of the explicit $p\times p$ interaction matrix $\mathbf{\bm{\Theta}}$. Problems with even higher values of $p$ are sometimes called ultrahigh-dimensional problems and require methods such as factorization machines or the penalized interaction estimation (PIE) technique proposed by Wang et al. (2021).
@@ Acknowledgments
This work made use of the High Performance Computing Resource in the Core Facility for Advanced Research Computing at Case Western Reserve University.
Research reported in this publication was supported by the National Library of Medicine of the National Institutes of Health under Award Number R01LM013311 as part of the NSF/NLM Generalizable Data Science Methods for Biomedical Research Program. The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health. This material is also based upon work supported by the National Science Foundation grant IIS-2318751.
@@ A.1 Targeted Interactions
Inclusion of the interaction terms between all the features might not be well suited given the conditions and the domain pertinent to a specific application. Instead, we may be interested in incorporating the interaction between a first set of features $\mathbf{\bm{X}}^{{}^{\prime}}\subseteq\mathbf{\bm{X}}$ with dimension of $r$ and a second set of features $\mathbf{\bm{X}}^{{}^{\prime\prime}}\subseteq\mathbf{\bm{X}}$ with dimension of $q$. We refer to these as targeted interactions, which may come from domain knowledge or specific interest in a set of interactions. With targeted interactions, the interaction matrix $\mathbf{\Theta}$ changes from a square $p\times p$ to a possibly rectangular $r\times q$} matrix representing the coefficients for the interaction terms between the features of $\mathbf{\bm{X}}^{{}^{\prime}},\mathbf{\bm{X}}^{{}^{\prime\prime}}$. We denote the latent representation matrices of the two sets of features $\mathbf{\bm{X}}^{{}^{\prime}},\mathbf{\bm{X}}^{{}^{\prime\prime}}$ as $\mathbf{\bm{Z}}^{{}^{\prime}}_{r\times d},\mathbf{\bm{Z}}^{{}^{\prime\prime}}_{q\times d}$, respectively.
@@ B.1 Linear Regression with Low Rank Model
Figure 6: Results of simulation experiments on linear regression for different values of the true latent dimension $d_{\text{true}}$. Lower RMSE denotes higher prediction accuracy. Results for $d_{\text{true}}=2$ are shown in Figure 2(a). Error bars denote one standard error.
@@
Figure 7: Results of simulation experiments on linear regression for different values of the chosen latent dimension $d$ and true latent dimension $d_{\text{true}}$. Lower RMSE denotes higher prediction accuracy. Error bars denote one standard error.
@@
Figure 8: Results of simulation experiments on logistic regression for different values of the true latent dimension $d_{\text{true}}$. Higher AUC denotes higher prediction accuracy. Results for $d_{\text{true}}=2$ are shown in Figure 2(b). Error bars denote one standard error.
@@
B.2 Logistic Regression with Latent Distance Model
@@
Finally, we explore what happens when we let $p$ grow to an extremely high value so that $\binom{p}{2}\gg n$. The results up to $p=500$, for which $\binom{p}{2}=124,\!750$ with n=1,000n=1,\!000 are shown in Figure 10. Notice that LIT-LVM continues to outperform elastic net until the point where all of the methods are barely better than a random guess at $p=500$.
B.3 LIT-LVM vs. Sparse Factorization Machines (FMs)
@@
In this simulation study, we evaluate the performance of sparse FMs compared to LIT-LVM through a series of classification experiments under varied conditions such as increased noise and sparsity injected in the interaction matrix $\mathbf{\bm{\Theta}}$. These conditions are critical for assessing the performance of the models, as they can potentially affect the predictive accuracy of models like FMs that rely on exact low-rank structures. By introducing greater noise variance and higher sparsity levels in the interaction weights during the simulations, we aim to demonstrate how each model handles deviations from the exact low-rank assumption, thereby providing insights into their practical use in scenarios where this assumption does not hold. Since we are interested in the accuracy of the FM and not the computation time, we formulate FMs as a special case of LIT-LVM with extremely large $\lambda_{l}=100,000$. (The model recovered by LIT-LVM is equivalent to that of an FM in the limit $\lambda_{l}\rightarrow\infty$ as discussed in Section 3.3.)
Figure 11 displays the mean AUC results for both models under baseline noise and sparsity conditions, while Figures 12 and 13 show the results when the models are subjected to increased noise levels and greater sparsity. Figure 12 shows the effect of increasing the noise variance to $4$ on the performance of the models, highlighting a significant drop in the accuracy of FMs due to their reliance on an exact low-rank structure of the interaction weight matrix. Similarly, Figure 13 demonstrates how increasing the sparsity level in the interaction weights impacts the performance of FMs compared to LIT-LVM. These experiments underscore the robustness of LIT-LVM under more challenging conditions where assumptions of the exact low-rank structures in FMs do not hold.
@@ C.1 OpenML Datasets
In our experiments, we utilize a diverse set of benchmark datasets to evaluate the performance of the proposed model. These datasets are chosen to ensure comprehensive coverage of different types of data distributions and characteristics. By using these datasets, we aim to thoroughly validate the performance and generalization ability of our model across both classification and regression tasks.
@@ C.1.1 Data Description
Table 4: Regression datasets from OpenML used for real data experiments, comprising a wide range of number of samples $n$ and features $p$. Datasets with a slash in the number of features $p$ have categorical variables, so the entry denotes before/after one-hot encoding. The ratio $p^{2}/n$ is computed using the number of features after one-hot encoding. The OpenML dataset ID column includes a link to the dataset.
Dataset Name # of Samples $n$ # of Features $p$ $p^{2}/n$ OpenML Dataset ID
boston 506 14/22 $0.957$ 531
bike_sharing_demand 1,737 6 $0.0207$ 44142
diamonds 53,940 10 $0.00185$ 44140
fri_c4_500_10 500 10 $0.200$ 604
medical_charges 163,065 12 $0.000883$ 44146
nyc-taxi-green-dec-2016 581,835 9 $0.000139$ 44143
pol 15,000 26 $0.0451$ 44134
rmftsa_ladata 508 10 $0.196$ 666
socmob 1,156 6/39 $1.32$ 541
space_ga 3,106 7 $0.158$ 507
superconduct 21,263 79 $0.294$ 44148
wind 6,574 15 $0.0342$ 503
@@
Dataset Name # of Samples $n$ # of Features $p$ $p^{2}/n$ OpenML Dataset ID
Bioresponse 3,434 420 $51.4$ 45019
clean1 476 169 $60.0$ 40665
clean2 6598 169 $4.33$ 40666
eye_movements 7,698 24 $0.0748$ 44157
fri_c4_500_100 500 101 $20.4$ 742
fri_c4_1000_100 1,000 101 $10.2$ 718
hill-valley 1,212 101 $8.41$ 1479
jannis 57,580 55 $0.0525$ 43977
jasmine 2,984 145 $7.05$ 45057
tecator 240 125 $65.1$ 851
We employ a range of datasets as detailed in Table 4 for regression and Table 5 for classification. These datasets are selected to test the model’s ability to handle different scenarios. The diversity in the number of samples and features ensures that our model is evaluated on a wide spectrum of regression and classification challenges. Links to these datasets are also provided for ease of access.
C.1.2 Hyperparameter Tuning
We use the hierNet R package implementing hierarchical lasso Bien et al. (2013). We tune its regularization parameter $\lambda$ over the same grid $\left[0.01,0.1,1,10,100\right]$ as for the other methods while leaving the $\ell_{2}$ penalty parameter fixed at $10^{-8}\lambda$, as the authors recommend not to tune this parameter. To match the other methods, we consider only interaction terms and exclude the “self-interactions” (entries on the diagonal of the interaction coefficient matrix).
C.1.3 Comparison of LIT-LVM with FMs
@@
Figure 15: Training and validation loss curves over epochs across different OpenML classification datasets. Each plot shows steady convergence, providing empirical evidence of stable optimization.
Figure 16: Prediction AUC on the clean2 dataset as $\lambda_{l}$ increases beyond $10^{5}$. The curve is nearly flat, indicating that LIT-LVM behaves like a sparse FM at large $\lambda_{l}$.
@@
C.2 Convergence of Optimization Procedure
To complement our empirical evaluations, we analyze the training dynamics of our model across various datasets. Specifically, we plot the training and validation loss curves over epochs in Figure 15, showing consistent and stable convergence of our optimization procedure.
C.3 Fitting Sparse Factorization Machine Models using LIT-LVM
To further validate that LIT-LVM with large $\lambda_{l}$ approximates the behavior of a sparse factorization machine (FM), we conduct an additional experiment on the clean2 dataset. Specifically, we increase $\lambda_{l}$ values beyond the $10^{5}$ value that we use in earlier experiments and measure its effect on prediction accuracy. As shown in Figure 16, the AUC remains nearly constant beyond this threshold, confirming that the LIT-LVM model with $\lambda_{l}=10^{5}$ is sufficiently large to represent a sparse FM model.
C.4 Survival Prediction and Compatibility Estimation for Kidney Transplantation
C.4.1 Data Preprocessing